how do you find a and b and c of a quadratic equation? I have data that is:
when x is 0, y is 1,.
when x is 1, y is 2.
when x is 2, y is 4.
when x is 3, y is 7.
when x is 4, y is 11.
when x is 5, y is 16.
and I have to write an equation for this data, which is quadratic I think. I don't know how to find a and b and c of the equation.

Answer by stanbon(75887)

how do you find a and b and c of a quadratic equation? I have data that is:
when x is 0, y is 1,.
when x is 1, y is 2.
when x is 2, y is 4.
when x is 3, y is 7.
when x is 4, y is 11.
when x is 5, y is 16.
and I have to write an equation for this data, which is quadratic I think. I don't know how to find a and b and c of the equation.
--------------
If it is quadratic, y = ax^2+bx+c and you need to find a,b, and c.
So,substitute x,y values from above to form 3 equations in a,b,c.
---------------------
(0,1) implies 1 = 0a+0b+c
(1,2) implies 2 = a + b + c
(2,4) implies 4 = 4a+2b+c
-----------------------------
Solve that system of equations by any means you know how.
I get a=(1/2), b=(1/2), c=1
=================
Cheers,
Stan H.